OpenStudy (anonymous):

How can you determine algebraically that the points A(1,2), B(2,4), C(3,6) are on the same line?

OpenStudy (amistre64):

\[\frac{Ay-By}{Ax-Bx}=\frac{Cy-By}{Cx-Bx}\]

OpenStudy (anonymous):

I'll use distance between points...

OpenStudy (amistre64):

distance doesnt tell you if they are in the same line; slope does

OpenStudy (anonymous):

let's see...

OpenStudy (amistre64):

three feet north is not in line with three feet west is it?

OpenStudy (anonymous):

AB=sqrt (5) bc=sqrt(5) AC=2swrt(5) so AC=AB+BC right?

OpenStudy (amistre64):

i got no idea what you typed

OpenStudy (anonymous):

distance between points A & B, as well B & C and A&C

OpenStudy (amistre64):

distance doesnt tell you if they are on the same line.

OpenStudy (amistre64):

equal distance could indicate a circle

OpenStudy (anonymous):

it has to in this case. how else AC would be = AB+BC...?

OpenStudy (amistre64):

oh isee what your saying now lol.... thats plausible then

OpenStudy (anonymous):

great!

OpenStudy (anonymous):

Nice job, can either of you take a look at the graphing question I posted a bit ago with the attachment? Thanks